OpenStudy (anonymous):

Annuities Problem. A man promised to pay $15000 at the end of each year. If money is 6% compounded monthly, what payment at the end of each month would be equally satisfactory?

OpenStudy (phi):

The future value of an "ordinary annuity" is \[ FV = payment \cdot \frac{(1+i)^n -1}{i} \] in your case, i = 0.06 and FV = 15000 I would use that info to solve for the payment

OpenStudy (anonymous):

what is n?

OpenStudy (anonymous):

i do not know of what is n?

OpenStudy (phi):

payment at the end of each month pay $15000 at the end of each year I would say 12, because after 12 months, 15000 is due

OpenStudy (anonymous):

So, by computation what's the final answer?

OpenStudy (phi):

are you saying you can't interpret the formula (i.e. use it ) ?

OpenStudy (anonymous):

I can but I need your answer if I have the right answer.

OpenStudy (phi):

post your work

OpenStudy (anonymous):

Answer $889.16

OpenStudy (anonymous):

right?

OpenStudy (phi):

yes, that looks good